OpenStudy (anonymous):

A flower garden is 4 feet longer than its width w. Write a polynomial that represents the area of the garden.

OpenStudy (anonymous):

Width=W Length= (W+4) Area= Length times Width so W(W+4)= W^2 + 4W hope this helps